A sample of N2 gas has a volume of 550 mL, has a pressure of 765 mm Hg, and is at a temperature of 27.8 °C. Calculate the amount (in moles) of N2 in the gas sample.
________ mol
1
Expert's answer
2019-03-21T06:25:01-0400

We shoud use Ideal Gas Law to answer this question:

"PV = nRT"

"765 mmHg\times0.550 L = n\times62.363\frac{L \times mmHg}{K \times mol}\times(28.7+273.15)K"

"n(N_2) =0.0224 mol"

Answer: 0.0224 mol